Academic Employment in Mathematics

AMS Standard Cover Sheet

This cover sheet is provided as an aid to departments in processing job applications. It should be included with your other application material. Please print or type. Do not send this form to the AMS.

Last (Family) Name:  \underline{Stein}
First Name or Initial:  \underline{William}
Middle Name or Initial: \underline{A}
Address through next June:  Cell Phone

Department of Mathematics

\underline{(617) 308-0144}

Harvard University e-mail address

Cambridge, MA 02138  was@math.harvard.edu

Current Institutional Affiliation  Web Page

Harvard University  modular.fas.harvard.edu

Highest Degree held or expected \underline{Ph.D.} Granting Institution \underline{University of California, Berkeley} Date (optional) \underline{May 2000} Ph.D. Advisor: \underline{Hendrik Lenstra} Ph.D. Thesis Title (optional) \underline{Explicit Approaches to Modular Abelian Varieties}

Indicate the mathematical subject areas in which you have done research using, if applicable, the Mathematics Subject Classification. If listing more than one number, list first the one number which best describes your current primary interest.

Primary Interest \underline{11} Secondary Interests (optional) \underline{11-04}

Give a very brief synopsis of your current research interests in the box below (e.g. finite group actions on four-manifolds). Avoid special mathematical symbols.

Elliptic curves, modular forms, L-functions, and explicit computation.

Most recent position held, if any, post Ph.D. University or Company Harvard University Position Title Benjamin Peirce Assistant Professor of Math. Dates \underline{2001-2005} Indicate the position for which you are applying and position posting code, if applicable Tenure or Tenure-Track If applying for a position which requires U.S. citizenship or U.S. permanent residency, indicate your eligibility. X Yes No If unsuccessful for this position, would you like to be considered for a temporary position?

Yes  X No  If yes, please check the appropriate boxes.
Postdoctoral Position  2+ Year Position  1 Year Position

List the names and affiliations of up to four individuals who will provide letters of recommendation if asked. X 1. \underline{Benedict Gross} X 2. \underline{Joe Harris} X 3. \underline{Hendrik Lenstra} X 4. \underline{Barry Mazur} X 5. \underline{Ken Ribet}